### v1

# Extracting the singularity of the logarithmic partition function

Abstract
In principle, the logarithm of the partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, in practice, this approach is often ineffective in numerical calculations, since resolving the singularities of the logarithmic partition function requires extremely high-precision data—especially when distinguishing between weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the leading volume-law term. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in $J\text{-}Q$ models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
Introduction.— Quantum phase transitions, spontaneous symmetry breaking, and the universality classes of critical phenomena are central concepts in condensed matter and statistical physics [1, 2, 3]. The type of phase transition is well defined by the singularity of free energy and its derivative. For a first-order phase transition, the first-order derivative of free energy is discontinuous. Otherwise, if its higher-order derivative is discontinuous, it is named a continuous phase transition. In principle, if the partition function can be calculated exactly, the phase transition will be fixed strictly according to the singularity. However, in practical numerical simulations, directly detecting such singularities through the free energy and its derivatives is difficult, as it requires extremely high-precision data. This challenge is particularly pronounced when distinguishing between weakly first-order phase transitions and deconfined quantum critical points  [4, 5, 6, 7, 8, 9, 10]. A major reason for this difficulty is that the singular contribution is typically overshadowed by a much larger, non-singular term that scales with the system volume. How to extract the singularity remains a significant challenge in theoretically computational physics.
Considering a quantum phase transition ($\beta\rightarrow\infty$), the singularity of free energy can be replaced by focusing on the logarithmic partition function. According to the finite-size scaling hypothesis [11, 12], an observable which is singular at $g_{c}$ (the parameter $g$ of Hamiltonian at the phase transition point) scales with the system size close to $g_{c}$ as a power of $L$ multiplied by a non-singular function of the ratio $\xi/L$, where $\xi$ is the correlation length and $L$ is the system size. Any singular quantity (not necessarily divergent) should thus be of the form:
$Q(\Delta g,L)=L^{\sigma}f(\Delta gL^{1/\nu})$ (1)
where $\Delta g=g-g_{c}$, $\nu$ is the critical exponent of the correlation length and $\sigma$ is an undetermined power. Since the singular part of the free energy is scale-invariant ($\sigma=0$[13, 11], the logarithmic partition function can be written in the following form:
$\ln Z(g)=a(g)L^{d}+f(\Delta gL^{1/\nu})$ (2)
where $Z$ is a partition function, $d$ is the spatial dimension and $a(g)L^{d}$ represents the leading term obeying a volume law. Here, we see that this dominant volume law term often masks the singular contribution of $f$, which explains the difficulty in numerically detecting the phase transition through such singularities.
An approach to extract the singular term $f$ is to construct a quantity in which the volume-law contributions cancel out, leaving only the singular part. Inspired by the method used to calculate topological entanglement entropy [14, 15], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\rightarrow\infty$) isolates the singular term $f$, with the dominant volume-law contributions canceled out. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)L^{d}+f$ and $2\ln Z(\beta)=2a(g)L^{d}+2f$. Coincidentally, we find that this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [16, 17, 18, 19, 20, 21, 22]:
$S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$ (3)
This definition can be generalized to an arbitrary order $n$, leading to the $n$-th order RTE: $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$. For simplicity, we focus on the 2nd-order RTE ( i.e., $n=2$) throughout this paper.
In this work, assisted by our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [23, 24, 25, 26, 27, 28, 29], we demonstrate the validity of our conjecture above: the RTE is indeed a dimensionless quantity that captures only the singular contribution. Furthermore, we observe a more pronounced singularity in the derivative of the RTE (DRTE) near the quantum critical point (QCP). We propose universal scaling relations for both the RTE and DRTE, which are confirmed by data collapse analyses performed near (2+1)D QCPs belonging to the O($N$) universality classes, with $1\leq N\leq 3$. More interestingly, since the leading non-singular term of $\ln Z$ is canceled out in the DRTE, it becomes a particularly powerful tool for detecting weakly first-order phase transitions even in normal sizes, which are notoriously difficult to resolve using conventional numerical approaches.
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. Additionally, we explore first-order transitions, with a particular focus on the challenging weakly first-order phase transitions (e.g., some controversial deconfined quantum critical points). Detailed descriptions of the model Hamiltonians will be provided in the following sections.
As shown by the definitions above, the 2nd-order Rényi thermal entropy can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 24] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 33, 34, 35]. Furthermore, following a procedure similar to that used in calculating the derivative of Rényi entanglement entropy and negativity [23, 36], the derivative of the RTE with respect to a general parameter $J$ can be written as:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
In this expression, the first average is evaluated on the ensemble corresponding to $Z(n\beta)$, while the second is based on $Z(\beta)$. In our simulations, the inverse temperature $\beta$ is set to be linearly proportional to the lattice size $L$ (i.e., $\beta=L$) in order to approach the quantum critical regime. Further methodological details are provided in the Supplementary Materials (SM).
Finite-size scaling form.— Considering a zero-temperature system and assuming the ground-state degeneracy is $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$+$f(\Delta gL^{1/\nu})$, where $f$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ can be used to probe both the degeneracy of discrete symmetry-breaking and the tower of excited states (TOS) associated with continuous symmetry-breaking.
More interestingly, as shown in Fig. 1, we observe that the curves for different system sizes cross at the critical point, which implies that $S^{(2)}$ is indeed a dimensionless quantity, consistent with our argument in the introduction. Near the critical point, the logarithmic partition function can be decomposed into a non-singular part $a(J)L^{d}$ and a dimensionless singular term $f(\Delta gL^{1/\nu})$. The non-singular contribution is the leading term, which makes it difficult to extract critical information directly from the partition function. As shown in Fig. 1, the crossings of curves for different sizes indicate that the RTE is independent of the system size. This provides strong numerical support for our argument that the leading volume-law contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel out in the definition of RTE, leaving only the dimensionless singular part. Additionally, we observe that the DRTE exhibits a stronger divergence near the QCP.
According to the finite-size scaling hypothesis  [11, 12], any singular quantity near a critical point should obey a scaling form as given in Eq.(1). In particular, the dimensionless RTE ($\sigma=0$), is expected to follow the scaling form:
$S^{(n)}(\Delta g,L)=f(\Delta gL^{1/\nu}).$ (5)
From Eq. (5), we can readily derive the scaling behavior of the DRTE:
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=f^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
where $f^{\prime}(x)$ is $\partial f(x)/\partial x$ and $x$ is a general parameter.
Figure 1: (a) The 2nd-order RTE and (c) its derivative, and (b) and (d) the data collapse analysis of the RTE and DRTE near the QCP of the bilayer Ising-Heisenberg model (with (2+1)D Ising universality class). We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Data of the dimerized anisotropic Heisenberg model with $\Delta=0.9$ (with (2+1)D O(2) universality class). We set $J_{2}=2.1035$, and take $J_{1}$ as the tuning parameter. (i–m) Data of the dimerized isotropic Heisenberg model with $\Delta=1$ (with (2+1)D O(3) universality class). We set $J_{2}=1$, and take $J_{1}$ as the tuning parameter. To achieve good data collapse, a subleading correction is also included: $Q(g,L)=L^{k}[f((g-g_{c})L^{1/\nu})+bL^{-1}]$.
In the following section, by expanding the scaling function $f(\Delta gL^{1/\nu})$ (or $f^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)d O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We consider the bilayer Ising-Heisenberg model with (2+1) D Ising transition  [37]. The Hamiltonian reads:
$H=-J_{z}\sum_{\alpha=1,2}\sum_{\langle ij\rangle}S_{i,\alpha}^{z}S_{j,\alpha}^{z}+J\sum_{i}{\mathbf{S}}_{i,1}\cdot{\mathbf{S}}_{i,2},$ (7)
in which $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ is the layer index. $-J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction. For $J/J_{z}\ll 1$, the ground state exhibits Ising order, breaking the $\mathbb{Z}_{2}$ symmetry. For $J/J_{z}\gg 1$, the ground state is in a spin-singlet states. The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [37].
We set $J=3.045$ and use $J_{z}$ as the tuning parameter. We then evaluate the 2nd-order RTE and its derivative in the vinicity of the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Fig. 1 (a) and (c). The RTEs of different sizes intersect at QCP, and their derivatives have a peak at QCP. We fit the scaling forms given by Eq. (5) and (6) to the data of RTE and its derivative separately. For the RTE data, we find $J_{z,c}=0.9999(1)$ and $\nu=0.62(1)$, while for the derivative of the DRTE data, we find $J_{z,c}=1.000(3)$ and $\nu=0.63(4)$. These results are consistent with the QCP $J_{z,c}=1$  [37] and the critical exponent $\nu=0.63012$ in the (2+1)D Ising universality class [38, 39, 37]. The universal scaling functions $f(\Delta gL^{1/\nu})$ and $f^{\prime}(\Delta gL^{1/\nu})$ obtained from the data collapse analysis are plotted in Fig. 1 (b) and (d).
Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [40, 41, 42]. The Hamiltonian reads:
$H=J_{1}\sum_{\langle ij\rangle}D_{ij}+J_{2}\sum_{\langle ij\rangle^{\prime}}D_{ij},$ (8)
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction and $\Delta$ is anisotropy parameter. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $10<\Delta<$ 1, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $10<\Delta<$ 1/ $\Delta>1$ / $\Delta=1$  [40, 41, 42].
The RTE and its derivative are evaluated around the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality). The results are shown in Fig. 1 (e,g) and (i,k), respectively. In each case, we fix $J_{2}$ and tune $J_{1}$ across the QCP, as described in the figure caption. Similar to the Ising transition case, we find that the RTE values for different lattice sizes intersect at the QCPs, and their derivatives exhibit peaks at the QCPs.
We fit the scaling forms given by Eq. (5) and (6) near the (2+1)D O(2) and O(3) QCPs separately. Around the O(2) QCP, we find $J_{2,c}=0.997(1)$ and $\nu=0.67(2)$ for the RTE data, and $J_{2,c}=0.999(1)$ and $\nu=0.66(2)$ for the DRTE data. These results are consistent with the known QCP and critical exponent: $J_{1,c}=1$ and $\nu=0.6703$ [38, 42].
Near the O(3) QCP, we find $J_{2,c}=0.523(1)$ and $\nu=0.70(1)$ for the RTE data, and $J_{2,c}=0.525(1)$ and $\nu=0.71(2)$ for the DRTE data. These results are also consistent with the known QCP and the critical exponent: $J_{1,c}=0.52337$ and $\nu=0.7073$ [38, 40, 41]. The universal scaling functions near O(2) and O(3) QCPs obtained from the data collapse analysis are shown in Fig. 1 (f), (h) and (g), (m) .
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$, and (b) and (c) the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model.
First-order and weakly first-order transitions.— Having established the universal scaling form for (2+1)D QCPs, we then investigate first-order transitions, especially the weakly first-order phase transition presenting significant challenges for traditional local physical quantities.
We first study the 2D checkerboard $J$-$Q$ model, which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [43]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, $2\partial\ln Z(\beta)/\partial J$ and $\partial\ln Z(2\beta)/\partial J$, as shown in Fig. 2 (b) and (c). The curve seems much more continuous compared with DRTE, because the derivative of logarithmic partition function includes a much larger term obeying volume-law. The phase transition of this model is obviously first-order according to the previous works [43, 23], thus we see that the cuves in (b) and (c) have very small bends around the phase transition point, especially for larger sizes.
The phase transition between AF Néel order and valence-bond solid in 2D quantum magnets can be (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [44, 45]. $J-Q_{2}$ and $J-Q_{3}$ models can realize the DQC in 2D quantum spin systems [46, 47]. This model has no sign problem and is, therefore, amenable to QMC simulations. Numerous QMC studies have shown that the phase transitions of the $J-Q$ model on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [46, 47, 48, 49, 50, 51]. However, the exponents of the $J-Q$ models do not satisfy the conformal bootstrap bounds of a $SO(5)$ critical point [52]. Recent studies have provided growing evidence suggesting that this DQC point is in fact a weakly first-order phase transition [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. In the following, we demonstrate that DRTE can clearly and simply prove that the phase transition in the $J-Q_{2}$ and $J-Q_{3}$ systems is a first-order phase transition.
The transition points of the two-dimensional square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [46, 47, 63, 62]. The numerical results of the DRTE for these two models are presented in Fig. 2 (d) and (g), where we also show the results for the derivative of logarithmic partition function itself in (e) (f) and (h) (i). The presence of a dominant background term makes it difficult for the derivative of the logarithmic partition function to directly reveal the weakly first-order nature of the phase transition. As shown in Fig. 2 (e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model, the curves appear smooth across the transition point, even for large system sizes. On the other hand, the DRTEs of both the $J-Q_{3}$ and $J-Q_{2}$ models exhibit clear discontinuities near the phase transition points. The results, by definition, prove that the phase transitions in both the $J-Q_{2}$ and $J-Q_{3}$ models are (weakly) first-order phase transitions. Notably, compared to $J-Q_{3}$ model, $J-Q_{2}$ model exhibits stronger finite-size effects, making it more difficult to determine its first-order nature prior to our study [62]. Our findings demonstrate that even weakly first-order phase transitions can be clearly detected using DRTE, even in normal sizes.
Conclusion and discussions.— We have proposed that the RTE only contains the singular term of logarithmic partition function, which is highly sensitive around phase transition points. Then we have systematically studied the RTE and its derivative around the QCPs in 2D quantum spin systems. We derived universal scaling forms for the RTE and DRTE, which capture the numerical results and yield the critical exponent $\nu$ through data collapse analysis near (2+1)D QCPs. More importantly, by definition, DRTE can clearly detect weak first-order phase transitions associated with deconfined quantum criticality, a task that is highly challenging for traditional local physical quantities.
Notably, RTE and DRTE is a universally applicable quantity because its definition does not depend on specific models or the precise definition of order parameters. Studying them has the potential to characterize phase transitions that are challenging for traditional local physical quantities, such as topological phase transitions, and to discover new universal critical exponents. Our work introduces a new and powerful paradigm for characterizing quantum phase transitions, including continuous phase transitions, first-order phase transitions, and the particularly challenging weakly first-order phase transitions.
Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Supplemental Materials
Appendix A Method
The $n$th order Rényi entropy is defined as $S^{(n)}=\frac{1}{1-n}\ln\mathrm{Tr}(\rho^{n})$ (or $\frac{1}{1-n}\ln\mathrm{Tr}(\rho_{A}^{n})$), where $\rho=e^{-\beta H}/Z$ and $Z=\mathrm{Tr}e^{-\beta H}$ (with $H$ being the Hamiltonian ). Here, $\rho_{A}=\mathrm{Tr}_{B}\rho$ represents the reduced density matrix of subsystem $A$ coupled with an environment $B$. The Rényi entanglement entropy is defined using $\rho_{A}$, while the Rényi thermal entropy (RTE) is based on $\rho$. In this work, we focus on RTE. From the above definitions, it is evident that the R’enyi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
The partition functions can be calculated using the bipartite reweight-annealing algorithm [23, 27, 64, 65] in conjunction with stochastic series expansion (SSE) QMC simulations [30, 12, 31, 32, 34, 33]. The key idea is to transform the problem of solving the ratio of partition functions with different parameter into the problem of reweighting:
$\frac{Z(J^{\prime})}{Z(J)}=\bigg{\langle}\frac{W(J^{\prime})}{W(J)}\bigg{\rangle}_{Z(\beta,J)orZ(n\beta,J)}$ (S1)
where the $\langle...\rangle_{Z(\beta,J)orZ(n\beta,J)}$ indicates that the QMC samplings is performed under the manifold $Z(\beta,J)orZ(n\beta,J)$ at parameter $J$ (note parameter $J$ is physical parameter here). $W(J^{\prime})$ and $W(J)$ represent the weights of a same configuration but under different parameters $J^{\prime}$ and $J$.
In this frame, we can designed the incremental process along the path composed of real physical parameters  [^1] It was virtual path before. In other words, all intermediates are physical partition functions at corresponding parameters. The continuously incremental trick here are introduce as follows:
$\frac{Z(J^{\prime})}{Z(J)}=\prod_{i=0}^{N-1}\frac{Z(J_{i+1})}{Z(J_{i})}$ (S2)
where $J_{N}=J^{\prime}$ and $J_{0}=J$, others $J_{i}$ locate between the two in sequence. In realistic simulation, we are able to gain any ratio ${Z(J^{\prime})}/{Z(J)}$ in this way. But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained. The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$. Considering a zero-temperature system and assuming the degeneracy of the ground state is $de$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln de$+$c((J-J_{c})L^{(1/\nu)})$, where $c$ represents the critical corrections, $J_{c}$ represents the critical point, and $\nu$ is the critical exponent of the correlation length. Therefore, we can use some states with known degeneracy as reference points, such as product states and discrete symmetry-breaking states.
The RTE derivative of $J$ is
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}\frac{dZ(n\beta)/dJ}{Z(n\beta)}-n\frac{dZ(\beta)/dJ}{Z(\beta)}\bigg{]}$ (S3)
According to $Z(\beta)=\mathrm{Tr}e^{-\beta H}$, we have
$\frac{dZ}{dJ}=\mathrm{Tr}[-\beta\frac{dH}{dJ}e^{-\beta H}]$ (S4)
Thus
$\frac{dZ/dJ}{Z}=-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}$ (S5)
Similarly, we can obtain
$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}\end{split}$ (S6)
Therefore, the RTE derivative can be rewritten as
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
The equations above is general and doesn’t depend on detailed quantum Monte Carlo (QMC) methods. Then let us discuss how to calculate them in stochastic series expansion (SSE) simulation. For convenience, we fix the Rényi index $n=2$ and choose the dimerized Heisenberg model in main text as the example for explaining technical details. The Hamiltonian is
$H=J_{1}\sum_{\langle ij\rangle}S_{i}S_{j}+J_{2}\sum_{\langle ij\rangle}S_{i}S_{j}$ (S8)
In the following, we fix $J_{2}=1$ and leave $J_{1}$ as the tunable parameter. Note the Hamiltonian is a linear function of $J_{1}$, that means $dH/dJ_{1}=H_{J_{1}}/J_{1}$ in which $H_{J_{1}}$ is the $J_{1}$ term in Hamiltonian. Then the RTE derivative can be simplified as
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}-2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S9)
In the SSE frame, it is easy to obtain $\langle H\rangle=\langle-n_{op}/\beta\rangle$  [^2] How to measure the energy in SSE has been carefully explained in Prof. Sandvik’s tutorial http://physics.bu.edu/~sandvik/programs/ssebasic/ssebasic.html, where $n_{op}$ is the number of the concerned Hamiltonian operators. Thus the Eq. (S9) can be further simplified to
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}-\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}+2\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S10)
where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones. It’s worth noting that there is no “$2$” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.