### v2

# Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions

Abstract
In principle, the logarithm of a partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, this approach is often ineffective in numerical calculations in practice, since resolving the singularity of a logarithmic partition function requires extremely high-precision data—especially when distinguishing weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the analytic terms. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in JQJ-Q models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
Introduction.— Quantum phase transitions, associated spontaneous symmetry breaking, and universality classes are central concepts in condensed matter and statistical physics [1, 2, 3]. The nature of a phase transition is determined by the singularity in the free energy or its derivatives: a first-order transition is marked by a discontinuity in the first derivative, while a continuous transition involves singular higher-order derivatives. In principle, an exact calculation of the partition function would allow one to identify the transition based on these singularities. However, in practice, detecting such features in numerical simulations is challenging, as it requires extremely high-precision data. This issue is particularly severe when distinguishing weakly first-order transitions from deconfined quantum critical points [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. A key difficulty arises because the singular contribution is often overwhelmed by a much larger, non-singular term that scales with the system volume. Extracting the singular part therefore remains a major challenge in computational physics.
For a quantum phase transition (inverse temperature $\beta\rightarrow\infty$; in the realistic simulations, the finite-size system gap $\Delta_{L}\sim\frac{1}{L^{z}}$, $\beta\sim L^{z}$ is usually used for probing the ground state properties, in which $L$ is the system size and $z$ is the dynamical critical exponent [17] ), the singularity of the free energy can be equivalently studied through the logarithm of the partition function. According to the finite-size scaling hypothesis [18, 17], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
$Q(\Delta g,L)=L^{\sigma}\tilde{S}(\Delta gL^{1/\nu})$ (1)
where $\Delta g=g-g_{c}$, $\nu$ is is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function. Since the singular part of the free energy is scale-invariant ($\sigma=0$[19, 18], the logarithm of the partition function can be expressed as:
$\ln Z(g,\beta,L)=a(g)\beta L^{d}+\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}(\Delta gL^{1/\nu})$ (2)
where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero or non-zero). The dominant volume-law term often obscures the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out. Inspired by the method used to extract topological entanglement entropy [20, 21], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\sim L^{z}$) can serve this role. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ and $2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$. Coincidentally, this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [22, 23, 24, 25, 26, 27, 28]:
$S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$ (3)
This definition generalizes naturally to arbitrary order $n$ : $S^{(n)}=\frac{1}{1-n}\ln[Z(n\beta)/Z(\beta)^{n}]$. For clarity and simplicity, we focus on the second-order RTE ($n=2$) throughout this work.
In this work, using our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [29, 30, 31, 32, 33, 34, 35], we validate the conjecture that the Rényi thermal entropy (RTE) is a dimensionless quantity capturing only the singular contribution. Moreover, its derivative (DRTE) exhibits a more pronounced singularity near the quantum critical point (QCP). We propose universal scaling forms for both RTE and DRTE, which are confirmed through data collapse analyses at (2+1)-dimensional (D) QCPs in the O($N$) universality classes, with $1\leq N\leq 3$. Notably, because the leading non-singular term in $\ln Z$ is canceled in DRTE, it provides a powerful diagnostic for weakly first-order phase transitions, even in relatively small systems (up to $L=56$), which are significant challenges to resolve using conventional numerical approaches.
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. We also examine first-order transitions, with a focus on weakly first-order cases, including some debated deconfined quantum critical points. The model Hamiltonians are described in detail in the following sections.
The second-order RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. We compute the partition functions using the bipartite reweight-annealing algorithm [29, 30] in combination with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 39, 40, 41]. Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [29, 42], the derivative of the RTE with respect to a general parameter $J$ is given by:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively. While the RTE is calculated using the reweight-annealing QMC method, the derivative (DRTE) is directly obtained from the two expectation values in Eq. (4); see Supplementary Materials (SM) for details. Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
Finite-size scaling form.—For a zero-temperature system with ground-state degeneracy $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$$\tilde{S}(\Delta gL^{1/\nu})$, where $\tilde{S}$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the tower of excited states (TOS) associated with continuous symmetry-breaking.
Near criticality, the logarithmic partition function consists of a dominant non-singular term $a(J)\beta L^{d}$ and a dimensionless singular term $\tilde{S}(\Delta gL^{1/\nu})$. The leading non-singular term obscures the critical behavior, making it difficult to extract the singular part directly from the partition function. More interestingly, as shown in Fig. 1, the curves of $S^{(2)}$ for different system sizes intersect at the critical point, indicates that the RTE is independent of system size, providing strong numerical evidence that the leading non-singular contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel exactly in the definition of RTE. This cancellation isolates the dimensionless singular term. Moreover, the DRTE displays a more pronounced divergence near the QCPs, further highlighting its sensitivity to critical fluctuations.
According to the finite-size scaling hypothesis  [18, 17], any singular quantity near a critical point follows a scaling form as given in Eq.(1). In particular, the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
$S^{(n)}(\Delta g,L)=\tilde{S}(\Delta gL^{1/\nu})$ (5)
From Eq. (5), we can readily derive the scaling behavior of the DRTE:
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
where $\tilde{S}^{\prime}(x)$ is $\partial\tilde{S}(x)/\partial x$ and $x$ is a general parameter.
Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality]. We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Results for the dimerized anisotropic Heisenberg model with $\Delta=0.9$ [(2+1)D O(2) universality], using $J_{2}=2.1035$ and tuning $J_{1}$. (i–m) Results for the dimerized isotropic Heisenberg model with $\Delta=1$ [(2+1)D O(3) universality], using $J_{2}=1$ and tuning $J_{1}$. To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$. $\omega$ is close to 1, and to reduce the numerical instability caused by increasing the number of fitting parameters, we set $\omega=1$. For $Q=S^{(2)}$, $k=0$. $k=1/\nu$ when $Q$ is DRTE.
In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)D O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [43]. The Hamiltonian is given by:
$H=-J_{z}\sum_{\alpha=1,2}\sum_{\langle ij\rangle}S_{i,\alpha}^{z}S_{j,\alpha}^{z}+J\sum_{i}{\mathbf{S}}_{i,1}\cdot{\mathbf{S}}_{i,2},$ (7)
where $\langle ij\rangle$ denotes the nearest-neighbor interaction, and $\alpha=1,2$ labels the two layers. $J_{z}<0$ represents the intralayer ferromagnetic (FM) Ising interaction, while $J>0$ represents the interlayer antiferromagnetic (AF) Heisenberg interaction. For $J/J_{z}\ll 1$, the ground state exhibits $\mathbb{Z}_{2}$ symmetry-breaking Ising order. In the opposite limit, $J/J_{z}\gg 1$, the system forms interlayer spin-singlet states. The QCP separating two phases is at $J/J_{z}=3.045(2)$ and belongs to the (2+1)D Ising universality class [43].
We set $J=3.045$ and use $J_{z}$ as the tuning parameter, studying the second-order RTE and DRTE near the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Figs. 1(a) and (c). The RTE curves for different system sizes intersect at the QCP, while the DRTE exhibits a clear peak at the critical point. We fit the scaling forms in Eqs. (5) and (6) to the RTE and DRTE data, respectively. The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.62(1)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE. These values are agreement with the known QCP $J_{z,c}=1$  [43] and the (2+1)D Ising universality class value $\nu=0.63012$  [44, 45, 43]. The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$, obtained from data collapse, are displayed in Figs. 1(b) and (d), respectively.
Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [46, 47, 48]. The Hamiltonian reads:
$H=J_{1}\sum_{\langle ij\rangle}D_{ij}+J_{2}\sum_{\langle ij\rangle^{\prime}}D_{ij},$ (8)
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction with anisotropy parameter $\Delta$. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $0<\Delta<1$, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [46, 47, 48].
We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs. 1(e,g) and (i,k), respectively. In each case, $J_{2}$ is fixed and $J_{1}$ is tuned across the transition, as detailed in the caption. Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points.
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. The inset shows data on the right side of the transition point. (b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model. The raw values of the partition function derivative vary significantly across system sizes (see SM), making direct comparison difficult. To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
We apply the scaling forms in Eqs. (5) and (6) to data near the (2+1)D O(2) and O(3) QCPs. For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.999(1)$, $\nu=0.66(2)$. These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [44, 48].
For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.70(1)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.71(2)$. These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [44, 46, 47]. The universal scaling functions near the O(2) and O(3) QCPs, obtained from data collapse, are shown in Figs. 1(f,h) and (j,m), respectively.
First-order and weakly first-order transitions.— Having established the universal scaling behavior at (2+1)D QCPs, we now turn to first-order phase transitions, with a focus on weakly first-order transitions, which presents significant challenges for traditional local physical quantities.
We first study the 2D checkerboard $J$-$Q$ model ( see SM for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [49]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, as shown in Figs. 2 (b) and (c). These curves appear much smoother across the transition compared with DRTE, due to the dominance of a non-singular, volume-law term. The phase transition of this model is obviously first-order according to the previous works [49, 29], thus we see that the curves in (b) and (c) have very small kinks around the phase transition point, especially for larger sizes.
The phase transition between AF Néel order and valence-bond solid (VBS) phase in 2D quantum magnets can appear (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [4, 5]. $J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [50, 51]. These models are free of the sign problem, making them ideal for QMC simulations. Extensive QMC studies have shown that the phase transitions of the JQJ-Q models on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [50, 51, 52, 53, 54, 55]. However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [56]. Recent works have increasingly suggested that this DQC scenario may instead correspond to a weakly first-order transition or as a nonunitary CFT  [7, 11, 57, 58, 12, 59, 60, 61, 8, 15, 62, 63, 64, 65, 16]. In the following, we show that the DRTE provides a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see SM for models’ details) are indeed first-order, even in rather small systems (up to $L=56$).
The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [50, 51, 66, 63]. The DRTE results for both models are shown in Figs. 2 (d) and (g), along with the derivatives of the logarithmic partition function in panels (e,f) and (h,i). Due to the dominant background contribution, the derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs. 2(e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model. This shows that the derivative of the logarithmic partition function can not be directly reveal the weakly first-order nature. In contrast, the DRTE reveals clear discontinuities at the transition points, providing direct evidence that both transitions are (weakly) first-order. Notably, the $J-Q_{2}$ model exhibits stronger finite-size effects than $J-Q_{3}$, which has made its first-order nature more difficult to resolve in previous studies [63]. Our results demonstrate that DRTE can unambiguously detect even weakly first-order transitions, in relatively small systems (up to $L=56$).
Conclusion and discussions.— We proposed that the RTE isolates the singular contribution of the logarithmic partition function, making it highly sensitive near phase transitions. We systematically studied the RTE and its derivative DRTE around QCPs in 2D quantum spin systems. By deriving universal scaling forms, we successfully captured the numerical data and extracted the critical exponent $\nu$ via data collapse near (2+1)D QCPs. Importantly, by definition, the DRTE provides a clear signature to distinguish weakly first-order phase transitions and deconfined quantum critical points even in small sizes, which is highly challenging to detect using conventional local observables.
Notably, RTE and DRTE are universally applicable, as their definitions do not depend on specific models or order parameters. They also have the potential to characterizing phase transitions that are difficult to probe with conventional local observables, such as topological transitions, and may reveal new universal critical behaviors. Our work establishes a general and effective framework for studying quantum phase transitions across a wide range of classes, including continuous, first-order, and particularly elusive weakly first-order transitions.
Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang and Wenan Guo. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Supplemental Materials
Appendix A Method
The $n$th order Rényi entropy is defined as $S^{(n)}=\frac{1}{1-n}\ln\mathrm{Tr}(\rho^{n})$ (or $\frac{1}{1-n}\ln\mathrm{Tr}(\rho_{A}^{n})$), where $\rho=e^{-\beta H}/Z$ and $Z=\mathrm{Tr}e^{-\beta H}$ (with $H$ being the Hamiltonian ). Here, $\rho_{A}=\mathrm{Tr}_{B}\rho$ represents the reduced density matrix of subsystem $A$ coupled with an environment $B$. The Rényi entanglement entropy is defined using $\rho_{A}$, while the Rényi thermal entropy (RTE) is based on $\rho$. In this work, we focus on RTE. From the above definitions, it is evident that the Rényi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
The partition functions can be calculated using the bipartite reweight-annealing algorithm [29, 33, 67, 68] in conjunction with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 40, 39]. The key idea is to transform the problem of solving the ratio of partition functions with different parameter into the problem of reweighting:
$\frac{Z(J^{\prime})}{Z(J)}=\bigg{\langle}\frac{W(J^{\prime})}{W(J)}\bigg{\rangle}_{Z(\beta,J)orZ(n\beta,J)}$ (S1)
where the $\langle...\rangle_{Z(\beta,J)orZ(n\beta,J)}$ indicates that the QMC samplings is performed under the manifold $Z(\beta,J)$ or $Z(n\beta,J)$ at parameter $J$ (note parameter $J$ is physical parameter here). $W(J^{\prime})$ and $W(J)$ represent the weights of a same configuration but under different parameters $J^{\prime}$ and $J$.
In this frame, we can designed the incremental process along the path composed of real physical parameters  [^1] It was virtual path before. In other words, all intermediates are physical partition functions at corresponding parameters. The continuously incremental trick here are introduce as follows:
$\frac{Z(J^{\prime})}{Z(J)}=\prod_{i=0}^{N-1}\frac{Z(J_{i+1})}{Z(J_{i})}$ (S2)
where $J_{N}=J^{\prime}$ and $J_{0}=J$, others $J_{i}$ locate between the two in sequence. In realistic simulation, we are able to gain any ratio ${Z(J^{\prime})}/{Z(J)}$ in this way. But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained. The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$: [Assuming that we have calculated the ratios $Z(n\beta,J^{\prime})/Z(n\beta,J)$ and $Z(\beta,J^{\prime})/Z(\beta,J)$ using the method described above, the ratio $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can be obtained if $Z(n\beta,J)/Z^{n}(\beta,J)$ is known. For a given zero-temperature phase, and assuming the ground state degeneracy is $d_{e}$, one can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z^{n}(\beta)}=\ln de$. Therefore, we can use states with known degeneracies — such as dimer product states (with degeneracy 1 in this work) and discrete symmetry-breaking states (with degeneracy 2 in this work) — as reference points at a given parameter $J$. In this way, $Z(n\beta,J)/Z^{n}(\beta,J)$ can be determined by solving: $\frac{1}{1-n}\ln\frac{Z(n\beta,J)}{Z^{n}(\beta,J)}=\ln de$.]
One might be concerned about how to deal with a Hamiltonian without a state with known degeneracy. In this way, we can introduce a temperature axis starting from infinite temperature and annealing down to the target temperature. The degeneracy at infinite temperature is known as $2^{N}$ (consider an N-spin system with spin-1/2).
The RTE derivative of $J$ is
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}\frac{dZ(n\beta)/dJ}{Z(n\beta)}-n\frac{dZ(\beta)/dJ}{Z(\beta)}\bigg{]}$ (S3)
According to $Z(\beta)=\mathrm{Tr}e^{-\beta H}$, we have
$\frac{dZ}{dJ}=\mathrm{Tr}[-\beta\frac{dH}{dJ}e^{-\beta H}]$ (S4)
Thus
$\frac{dZ/dJ}{Z}=-\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}$ (S5)
Similarly, we can obtain
$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}\end{split}$ (S6)
Therefore, the RTE derivative can be rewritten as
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
The equations above is general and doesn’t depend on detailed quantum Monte Carlo (QMC) methods. Then let us discuss how to calculate them in stochastic series expansion (SSE) simulation. For convenience, we fix the Rényi index $n=2$ and choose the dimerized Heisenberg model in main text as the example for explaining technical details. The Hamiltonian is
$H=J_{1}\sum_{\langle ij\rangle}S_{i}S_{j}+J_{2}\sum_{\langle ij\rangle}S_{i}S_{j}$ (S8)
In the following, we fix $J_{2}=1$ and leave $J_{1}$ as the tunable parameter. Note the Hamiltonian is a linear function of $J_{1}$, that means $dH/dJ_{1}=H_{J_{1}}/J_{1}$ in which $H_{J_{1}}$ is the $J_{1}$ term in Hamiltonian. Then the RTE derivative can be simplified as
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}-2\beta\bigg{\langle}\frac{H_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S9)
In the SSE frame, it is easy to obtain $\langle H\rangle=\langle-n_{op}/\beta\rangle$  [^2] How to measure the energy in SSE has been carefully explained in Prof. Sandvik’s tutorial http://physics.bu.edu/~sandvik/programs/ssebasic/ssebasic.html, where $n_{op}$ is the number of the concerned Hamiltonian operators. Thus the Eq. (S9) can be further simplified to
$\frac{dS^{(2)}}{dJ_{1}}=\bigg{[}-\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z(n\beta)}+2\bigg{\langle}\frac{n_{J_{1}}}{J_{1}}\bigg{\rangle}_{Z}\bigg{]}$ (S10)
where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones. It’s worth noting that there is no “22” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.
Note that the RTE is computed using the reweight-annealing QMC method mentioned above, while the DRTE is obtained directly by separately evaluating the two expectation values involved in Eq. (S9) or Eq. (S10). This implies that the computational cost of evaluating DRTE is roughly twice that of energy calculations. Typically, we perform on the order of one million Monte Carlo steps when computing the energy, and we have adopted the same number of steps in the present study.
Figure S1: $J$-$Q$ models, in which the antiferromagnetic Heisenberg interaction ($J$-term) acts on all nearest-neighbor bonds. (a) The checkerboard $J$-$Q$ model on a square lattice with periodic boundary condition in both directions. The four-spin interaction ($Q$-term) acts on nearest-neighbor bond pairs that form a plaquette. (b) $J-Q_{2}$ model: the four-spin $Q$ interaction covers the entire lattice. (c) $J-Q_{3}$ model: the six-spin $Q$ interaction covers the entire lattice.
Appendix B JQJ-Q models
In this work, we studied three types of $J$-$Q$ models, namely checkerboard $J$-$Q$, $J$-$Q_{2}$, and $J$-$Q_{3}$ models. The Hamiltonian of the checkerboard $J$-$Q$ model on a square lattice is given by [49]
$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijkl]}P_{ij}P_{kl},$ (S11)
[…]
$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijklmn]}P_{ij}P_{kl}P_{mn},$ (S12)
, as illustrated in Fig. S1(c). The DRTE results for three models are shown in Figs. S2, along with the derivatives of the logarithmic partition function in panels (b,c), (e,f) and (h,i).
Figure S2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. An inset shows the data on the right side of the phase transition point. (b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model.